# Efficient Transfer Learning for Video-language Foundation Models

### v1

@@ 4 Experiments
4.1 Datasets
Kinetics-400 [1] is a large-scale video dataset containing approximately 240,000 training videos and 20,000 validation videos across 400 human action categories, with an average video length of 10 seconds. Due to its high quality, it has become one of the most popular benchmarks for video recognition.
Kinetics-600 [2] is an extension of Kinetics-400, with approximately 392,000 training videos, 30,000 validation videos, and 60,000 test videos in 600 human action categories. It introduces 220 new action categories not present in Kinetics-400. We evaluate the zero-shot performance on these 220 new categories and use three splits provided by previous work [4]. The test set is used for evaluation, and we report the average performance across the three splits.
UCF-101 [31]is an action recognition dataset containing 13,320 videos in 101 action categories, collected from YouTube. It has three official splits for training and validation data.
HMDB-51 [20] contains 7,000 videos across 51 action categories, collected from movie clips and web videos. The dataset has three official splits, with 3,570 training videos and 1,530 validation videos in each split.
Something-Something V2 [13] is a temporally demanding dataset that requires fine-grained temporal understanding. It includes 220,000 videos across 174 action categories.
@@ 4.2 Experimental Setup
Architecture. We employ the ViCLIP [36] pre-trained ViT-B/16 in our experiments. For the zero-shot transfer, few-shot learning, base-to-novel generalization and fully-supervised settings, we introduced the proposed multi-modal spatio-temporal adapter starting from the $k=1/8/1/1$-th transformer block to the last one in both the language and vision branches, respectively.
@@ 4.4 Ablation Studies
Scaling Factor $\lambda$. The scaling factor $\lambda$ is a critical hyperparameter that controls the balance between general and task-specific features within the MSTA. As shown in Table 6(c), we thoroughly evaluated various scaling factors. Our results reveal that a scaling factor of $\lambda=0.005$ provides the best performance trade-off (HM) between base and novel categories. This optimal balance is essential for enabling the model to quickly adapt to base categories without sacrificing its ability to generalize to novel categories. In contrast, larger scaling factors, while improving performance on base categories, can hinder the model’s performance on novel categories. Smaller scaling factors, on the other hand, can limit the model’s adaptability to downstream tasks.
Weighting Factor $\alpha$. The weighting factor $\alpha$ associated with the consistency constraint loss is a key hyperparameter that influences model performance. As shown in Table 6(d), the choice of $\alpha$ has a significant impact on accuracy. The model achieves peak performance when $\alpha=1.0$. Deviations from this value, whether by increasing or decreasing $\alpha$, lead to a decline in performance. This finding underscores the importance of carefully tuning the consistency constraint to maintain a balance between adaptability and generalization capabilities.
Variants of Adding MSTA. We evaluate the impact of adding MSTA to different encoder layers, with results summarized in Table 6(f). The performance of MSTA improves consistently as the number of added layers increases. Furthermore, given the same number of layers, MSTA performs better when placed higher in the network (i.e., farther from the input). For example, inserting MSTA into layers 7 to 12 yields 0.3 higher HM than inserting into layers 1 to 6, despite both configurations using six MSTA layers. Notably, the insertion of MSTA in layers 8 to 12 delivers the best performance for base class tasks, which is particularly advantageous for few-shot learning scenarios where rapid adaptation to new categories is essential.
Impact of different descriptions. We investigate the impact of spatiotemporal descriptions on the performance of our proposed method. The results presented in Table 7 show that each description complements the others, highlighting the importance of both spatiotemporal descriptions for recognition tasks. Additionally, we observe that the temporal descriptions have a more pronounced effect on performance compared to the spatiotemporal descriptions.
Table 7: Study on different descriptions.
Description Base Novel H
w/o DES 68.2 53.2 59.77
w/o Spatio 68.4 53.4 59.97
w/o Temporal 68.8 53.3 60.06
Full 68.6 53.5 60.12
Cost Analysis. We analyze the additional costs of our method during training and inference in Table 8. Latency is measured in our baseline training setup, while throughput is evaluated using the largest batch size that can be processed without exceeding the memory limit of a single NVIDIA A100-80G. Our pipeline incurs only an additional 0.1× training time and results in a throughput reduction of approximately 5%, which is acceptable considering the performance gains achieved.
Table 8: Additional cost analysis of our method, we report step latency during training, and throughput (TP) during inference. We refer to Top-1 as zero-shot accuracy on Kinetics-600.
Description Top-1 (%) Latency (s) TP (video/s)
ViCLIP [36] 71.5 0.45 (1.0x) 69.3(1.0x)
Ours 74.5 0.51 (1.1x) 65.7 (0.95x)
@@ 5 Conclusion
In conclusion, this paper presents a novel Multi-modal Spatio-Temporal Adapter (MSTA) for efficient transfer learning in video-language foundation models. The MSTA is designed to enhance the alignment between text and video representations, thereby achieving a balance between general knowledge and task-specific knowledge. Our method introduces a spatio-temporal description-guided consistency constraint to mitigate over-fitting and enhance generalizability, which is particularly beneficial in data limited scenarios. Through extensive experiments across various tasks, our MSTA demonstrates outstanding performance and achieves state-of-the-art results while using a minimal number of trainable parameters. The comprehensive ablation studies further validate the effectiveness and scalability of our proposed method. Overall, our work contributes significantly to the field of video action recognition by providing an efficient and effective transfer learning approach that preserves the generalizability of pre-trained models while adapting them to specific downstream tasks.
6 Limitation and Broader Impact
Limitation. Although our method achieves excellent results across various video recognition settings, there are still areas for improvement. For example, MSTA connects the gradients of visual and textual features. Despite using gradient scaling to prevent instability, we believe that the backpropagation performed under this setup is not optimal. Therefore, exploring better ways to train this type of module could further enhance MSTA’s performance. On the other hand, using spatiotemporal descriptions generated by LLMs for constraints may lead to performance degradation due to the lower quality of these descriptions. In future work, we will further investigate how to leverage LLMs to improve the efficient transfer of video-language models, aiming to explore the generality and versatility of our approach.
Broader Impact. The adaptability of foundational models to various downstream tasks has become a major focus in the current field of machine learning [17, 5, 25]. We believe that retaining the existing capabilities of foundational models while integrating new knowledge is an important issue worth exploring in depth. We hope this research provides new insights and approaches for the broader and long-term application of foundational models. Our work focuses on video recognition tasks, which have widespread applications in real-world scenarios such as surveillance. However, before these technologies can be deployed, it is crucial to carefully address potential issues related to privacy protection and rights.
7 Pseudo-code
The pseudo-code implementation of MSTA is provided below, along with a concise definition of MSTA and its insertion method.
Algorithm 1 Implementation of Text Adapter & Shared Adapter in PyTorch-like style
  def _build_adapter(d_model, n_layers, l_start, l_end, mid_dim, dropout_rate=0.1):
        adapter = [None] * (n_layers + 1)
        for i in range(l_start, l_end+1):
            if mid_dim == d_model:
                adapter[i] = nn.Sequential(
                    nn.Linear(d_model, mid_dim),
                    nn.ReLU()
            else:
                adapter[i] = nn.Sequential(OrderedDict([
                    (”down”, nn.Sequential(
                    nn.Linear(d_model, mid_dim),
                    nn.ReLU()
                    (”dropout”, nn.Dropout(p=dropout_rate)),
                    (”up”, nn.Linear(mid_dim, d_model))
        adapter = nn.ModuleList([a for a in adapter])
        return adapter
Algorithm 2 Implementation of Visual Adapter in PyTorch-like style
  def _build_visual_adapter(d_model, n_layers, l_start, l_end, mid_dim, dropout_rate=0.1):
        adapter = [None] * (n_layers + 1)
            for i in range(l_start, l_end+1):
                adapter[i] = nn.Sequential(OrderedDict([
                (”down”, nn.Sequential(
                nn.Linear(d_model, mid_dim),
                nn.ReLU()
                (”dropout”, nn.Dropout(p=dropout_rate)),
                (”up”, nn.Linear(mid_dim, d_model)),
                (”temporal_up”, nn.Conv3d(mid_dim, d_model, kernel_size=(3,1,1), stride=(1, 1, 1),
                padding=”same”, groups=mid_dim, dilation=(1, 1, 1))),
        adapter = nn.ModuleList([a for a in adapter])
        return adapter

### v4

@@ 4.1 Experimental Setup
Dataset. We conducted experiments on six video benchmarks: Kinetics-400 [1], Kinetics-600 [2], UCF-101 [31], HMDB-51 [20], and SomethingSomething V2 [13]. Our study covers multiple settings, including zero-shot transfer, few-shot learning, base-to-novel generalization and fully-supervised video recognition.
Architecture. We employ the ViCLIP [36] pre-trained ViT-B/16 in our experiments. For the zero-shot transfer, few-shot learning, base-to-novel generalization and fully-supervised settings, we incorporated our multi-modal spatio-temporal adapter starting from the 1st, 8th, 1st, and 1st transformer blocks and extended it to the last block in both the language and vision components.
@@ 4.3 Ablation Studies
Scaling Factor $\lambda$. The scaling factor $\lambda$ balances general and task-specific features in MSTA. As shown in Table 6(c), we found $\lambda=0.005$ achieves the best trade-off (HM) between base and novel categories. Larger values enhance base-category performance but reduce generalization, while smaller values limit adaptability.
Weighting Factor $\alpha$. The weighting factor $\alpha$, associated with the consistency constraint loss, is a critical hyperparameter affecting model performance. As demonstrated in Table 6(d), varying $\alpha$ significantly impacts model accuracy, with optimal performance observed at $\alpha=1.0$. Deviating from this optimal value, either by increasing or decreasing $\alpha$, results in performance degradation. These observations highlight the necessity of precise tuning of the consistency constraint to achieve an effective balance between model adaptability and generalization.
Numbers of Descriptions. We investigate the influence of varying the number of descriptions in Table 6(e). We conducted experiments with 2, 4, and 8 spatio-temporal descriptions. It can be observed that the performance is better when $N=2$. This may be because as $N$ increases, the hallucination problem of the LLM becomes more severe, resulting in a significant amount of noisy descriptions.
Variants of Adding MSTA. We investigate the impact of incorporating MSTA into various encoder layers, with the results summarized in Table 6(f). The performance consistently improves as the number of MSTA layers increases. Additionally, for the same number of MSTA layers, placing them in higher encoder layers (farther from the input) yields superior results. For instance, integrating MSTA into layers 7 to 12 achieves an HM score 0.3 points higher than integrating it into layers 1 to 6, despite both configurations involving six MSTA layers. Importantly, inserting MSTA into layers 8 to 12 provides the best performance for base class tasks, highlighting its particular benefit in few-shot learning scenarios that demand rapid adaptation to novel categories.
@@ 5 Conclusion
In conclusion, this paper introduces a novel Multi-modal Spatio-Temporal Adapter (MSTA) designed for efficient transfer learning in video-language foundation models. The MSTA aims to enhance alignment between text and video representations, achieving an optimal balance between general knowledge and task-specific knowledge. The proposed method incorporates a spatio-temporal description-guided consistency constraint to reduce overfitting and improve generalization, particularly beneficial in scenarios with limited data. Extensive experiments conducted on diverse tasks demonstrate outstanding performance and state-of-the-art results, achieved with minimal trainable parameters. Comprehensive ablation studies further confirm the effectiveness and scalability of the proposed approach. Overall, our work significantly advances the field of video action recognition by offering an efficient transfer learning framework that maintains the generalizability of pre-trained models while effectively adapting them to specific downstream tasks.